OpenStudy (anonymous):

A pole that is 20 feet tall casts a shadow 48 feet long. At the same time, another pole casts a shadow having a length of 12 feet. What is the height, in feet, of the second pole? Answers: a. 5 b. 8.5 c. 28.8 d. 80 help please :3

OpenStudy (anonymous):

(48/20)x12=28.8

OpenStudy (anonymous):

You can solve this by setting up proportions: \[20/48=12/x\] Just make sure you keep the length of the shadow/ length of the pole in the same place in each fraction. Now you just have to solve for x.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

By doing 48 divided by 20 you find the length of the shadow for every feet then by multiplying by 12 you get it for 12 feet